# An Accurate Efficient Analytic Model of Fidelity under Depolarizing Noise oriented to Large Scale Quantum System Design

### v1

@@ Abstract
Fidelity is one of the most valuable and commonly used metrics for assessing the performance of quantum circuits on error-prone quantum processors. Several approaches have been proposed to estimate circuit fidelity without the need of executing it on quantum hardware, but they often face limitations in scalability or accuracy. In this work, we present a comprehensive theoretical framework to predict the fidelity of quantum circuits under depolarizing noise. Building on theoretical results, we propose an efficient fidelity estimation algorithm based on device calibration data. The method is thoroughly validated through simulation and execution on real hardware, demonstrating improved accuracy compared to state-of-the-art alternatives. The proposed approach provides a scalable and practical tool for benchmarking quantum hardware, comparing quantum software techniques such as compilation methods, obtaining computation bounds for quantum systems, and guiding hardware design decisions, making it a critical resource for the development and evaluation of quantum computing technologies.
keywords:
Quantum Information, Quantum Computing, Depolarizing Noise, Quantum Fidelity
@@ 1 Introduction
Quantum fidelity [20, 21] is a measure of the similarity between two quantum states and hence serves as an indicator of either the reliability of a computation or the presence of excessive errors. It is calculated using the density matrices representing the quantum states ($\rho$ and $\sigma$), as defined in Equation (1).
@@
Estimating the fidelity of quantum circuits without executing them on quantum hardware is crucial to advance in the design of quantum processors. A reliable fidelity estimation method enables researchers to compare quantum systems, optimize compilation strategies, and predict performance across different hardware platforms. As the number of qubits increases, the estimation method must remain computationally efficient to scale with the growing complexity of quantum circuits.
@@
Simulating quantum circuits [22, 23, 24, 25] enables adding noise to the quantum state, a process known as statistical fault injection [26]. The fidelity of the computation can be assessed by comparing a simulated noisy state with its noiseless counterpart. However, the simulation of quantum states is computationally expensive, with complexity growing exponentially with the number of qubits, restricting its application to small-scale circuits [27]. The accuracy of the estimation will depend on the noise model used in the simulation, as well as on the precision of the simulation itself, which may introduce errors arising from approximations (e.g., limiting the dimensionality of tensors in a tensor network simulation [28]).
Analytical approaches for fidelity estimation [29, 30] typically rely on gate error rates obtained from device calibration, usually through randomized benchmarking [31, 32], without assuming any error model in particular. An example of such methods is the Probability of Successful Trial (PST) [29], also known as Estimated Success Probability (ESP), a commonly used metric for estimating the fidelity of a quantum computation by multiplying the individual gate fidelities ($F_{g_{i}}$) and measurement fidelities ($F_{m_{i}}$) of the circuit (Equation (4)). However, ESP does not account for how errors propagate through the system or specify which qubits are affected by the errors, obtaining a circuit fidelity that does not depend on the circuit structure but just on the number of gates and their respective fidelities.
@@
The Quantum Vulnerability Analysis (QVA) method [30], like ESP, uses individual gate fidelities as its basis. However, it introduces a hyperparameter ($w\in[0,1]$) that quantifies the degree of cross-error introduced by a two-qubit gate, requiring fine-tuning for each specific system. Additionally, unlike ESP, QVA accounts for the fidelity of each two-qubit gate twice (once for each qubit involved), resulting in consistently lower fidelity estimates compared to ESP. Machine-learning-based techniques [33, 34, 35] have also been proposed to estimate circuit fidelity. These approaches often target specific platforms, requiring extensive training data and retraining for different devices.
To better understand the strengths and limitations of existing fidelity estimation methods, we compare their predictions against experimental results from a real quantum processor.
Figure 1 compares the fidelity predictions of statistical fault injection (using Qiskit’s simulator [36]), ESP, and QVA for quantum circuits ranging from 2 to 8 qubits, against the success rate of the same circuits executed on IBM Q Kyiv (Eagle r3 processor [37]). The process from obtaining the circuits to computing the success rate after readout mitigation is explained in Section 4.2. Figure 1 shows how both Qiskit’s simulation and ESP inaccurately predict the fidelity observed in real quantum processors, consistently overestimating it. Conversely, QVA’s performance highly depends on its hyperparameter ($w$), with fidelity predictions varying significantly, up to 0.8, between the extreme values of $w=0$ and $w=1$ (QVA max and QVA min in Figure 1). This introduces the additional challenge of determining the optimal $w$, which the original work [30] addressed by training a machine learning model for this purpose. In our analysis, we use $w=0.5$ (depicted as red points in Figure 1) as it provides a balanced prediction (see box plot in Figure 1), while also reporting results for $w=0$ and $w=1$. In the original work [30] the authors identify $w=0$ as the optimal value for large circuits.
@@
These observations highlight the need for a more accurate fidelity estimation model that, like QVA and ESP, is scalable in terms of computational cost, allowing to increase the number of qubits and gates in the circuit. However, it is essential to minimize the dependence on hyperparameters that can excessively influence predictions, as this variability can compromise the reliability and consistency of the model’s outcomes.
In this work, we develop a theoretical model for estimating the fidelity under depolarizing noise (Equation (5)), in which each quantum gate introduces depolarization to the involved qubits, a common assumption regarding operational errors [38, 39, 40, 41, 42, 2]. By examining how depolarizing noise impacts circuit fidelity, we lay the groundwork for a scalable fidelity estimation algorithm rooted in theoretical principles.
@@
We will primarily use the last expression in Equation (5) to represent the depolarizing channel, where $d$ is the dimensionality of the quantum system being depolarized, $I_{d}/d$ is the maximally mixed state with the same dimensions ($d$) as $\rho$ (the quantum state described by its density matrix), and $p$ is the depolarization factor (or depolarization probability). This formulation provides a compact and general framework to model depolarization effects across systems of varying dimensions.
2 Results
2.1 Theoretical Modeling of the Fidelity Loss
@@
To further explore the impact of depolarization, Corollary 1 analyses the fidelity loss after a single depolarization, considering the effects of repeated depolarizations over time. This result will be used to quantify each gate error’s contribution when estimating the fidelity of quantum circuits.
@@
Figure 2 presents a comparison between the fidelities predicted by our theoretical model and those obtained from statistical fault injection simulations. The results span 93 quantum circuits, varying in size from 2 to 8 qubits and containing up to 5000 gates. Circuits are compiled to match with the restrictions of the superconducting processor used in latter experiments (see Section 4.2).
@@
As expected, the fidelity predictions align closely with the simulated values, though slight discrepancies arise due to the theoretical upper and lower bounds employed in our model. To better illustrate these discrepancies, Figure 2 also shows the absolute difference between the predicted (with $p_{ent}=0.5$) and simulated fidelities. The observed deviations remain within a reasonable range (for all data points the absoulte fidelity difference remains under 0.07), confirming that our approach provides a robust approximation of circuit fidelity while being significantly more efficient than direct simulation.
@@
Figure 2: Comparison of predicted versus simulated fidelities with single-qubit gates depolarization of $p_{1}=10^{-3}$ and two-qubit gates depolarization of $p_{2}=5\cdot 10^{-3}$. In the left plot, each point illustrates the predicted ($x$-axis) against the simulated fidelity ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. Horizontal lines indicate the range of predicted fidelities arising from the variability of the hyperparameter $p_{ent}$. The right plot shows the difference on fidelity ($fid_{pred}-fid_{simul}$).
2.2 Quantum Circuit Fidelity Estimation
Having validated the model through simulation, we now assess its performance against real quantum hardware. Using the same set of compiled circuits, we compare the predicted fidelities to the actual success rates (i.e., the number of correct measurements divided by the total number of executions) measured on an IBM Q superconducting processor [37].
The fidelity was estimated using our model alongside other methods, including Qiskit simulation (using noise models associated with the used processor), ESP, and QVA. We compared these predictions to the success rates obtained from executions on quantum hardware after readout error mitigation. Additionally, a version of each model incorporating coherence errors ($T_{1}$ and $T_{2}$), detailed in Section 4.1, were also considered to provide a more comprehensive assessment of a real execution. Since quantum hardware is subject to additional sources of error beyond those captured by the considered models, it is not possible to achieve an accurate prediction by considering only gate errors, making the inclusion of $T_{1}$ and $T_{2}$ essential for a robust validation against quantum hardware.
The execution flow and the detailed implementation of the proposed fidelity estimation algorithm are presented in Section 4.
Figure 3 illustrates the measured success rate in the quantum device (IBM Q Eagle processor) after readout error mitigation (on the $y$-axis) versus the predicted fidelity (on the $x$-axis) for each estimation method. In this plot, a perfect prediction would place the data point along the diagonal. For the QVA and proposed models, horizontal lines indicate the influence of the hyperparameters in each algorithm (i.e., $w$ for QVA and $p_{ent}$ for the proposed model), with the reported point representing the average value between the extremes of these parameters. Most gate-error models tend to estimate a higher fidelity than the actual observation on the real device, which is expected since the models do not account for all types of errors present in actual hardware (e.g., coherence or crosstalk errors). This effect is evident when incorporating coherence errors into the estimation models, as it demonstrates a notable improvement in aligning predictions with the observed success rates for all models except for QVA, where it leads to a worse prediction, estimating a lower fidelity than the observed one.
Figure 3: Comparison of predicted versus measured fidelities for all estimation models. Each scatter plot illustrates the predicted fidelity ($x$-axis) against the measured success rate ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. For the QVA and the proposed model, horizontal lines indicate the range of predicted fidelities arising from the variability of their hyperparameters $w$ and $p_{ent}$ respectively.
Interestingly, QVA sometimes predicts a lower fidelity than the observed success rate despite not including coherence or crosstalk errors in its model. This discrepancy arises because QVA decreases the global fidelity twice for each two-qubit gate, doubling the reduction in fidelity compared to the actual gate error. When coherence errors are incorporated into the QVA model, the predicted fidelities decrease even further, as shown in the lower row of Figure 3. Moreover, the long horizontal lines for the QVA in Figure 3 highlight the significant influence that the hyperparameter for the cross-error ($w$) has on its predictions.
Across all estimation techniques, we observe that, when only taking into account gate errors (i.e., upper row), for circuits with either very few or very many gates, the predictions tend to be relatively accurate. The explanation lies in the simplicity of estimating fidelity in these cases. For circuits with a very low number of gates, the fidelity is typically close to one, and fewer interactions among qubits reduce cross-talk errors, which are not considered by none of the evaluated models. Additionally, coherence noise effects (which are not accounted for in ESP and QVA) are less significant due to the shorter execution times of such circuits. Conversely, for circuits with a high number of gates, the fidelity generally approaches zero, causing the absolute difference between the predicted and real success rates to decrease.
To assess the accuracy of the proposed prediction models, we compute several regression metrics typically used to evaluate the performance of predictive models. These include the Mean Absolute Error (MAE) [43], which measures the average absolute difference between predicted and actual values; the Mean Squared Error (MSE) [44], which measures the average of the squares of the errors, that is, the average squared difference between the estimated values and the actual value; and the $R^{2}$ Score [45], which quantifies the proportion of variance in the observed data that is predictable from the model. Finally, the Pearson Correlation [46] coefficient is used to assess the linear relationship between the predicted and actual values, where a higher value indicates a stronger linear relationship. Additionally, the results with the cross-error parameter set to zero ($w=0$) in the QVA algorithm are also included since, in the analysis shown in [30], for most large circuits (over 100 two-qubit gates), the optimal $w$ was close to 0.
Mean Absolute Mean Squared $R^{2}$ Score $\uparrow$ Pearson
Error (MAE) $\downarrow$ Error (MSE) $\downarrow$ Correlation $\uparrow$
Qiskit 0.422 0.273 -0.831 0.606
Qiskit +$T_{1,2}$ 0.094 0.024 0.843 0.927
ESP 0.229 0.098 0.341 0.827
ESP +$T_{1,2}$ 0.080 0.016 0.891 0.945
QVA 0.134 0.034 0.769 0.903
QVA +$T_{1,2}$ 0.161 0.051 0.658 0.898
QVAw=0 0.126 0.043 0.715 0.895
QVAw=0 +$T_{1,2}$ 0.088 0.015 0.899 0.951
This Work 0.210 0.084 0.434 0.849
This Work +$T_{1,2}$ 0.067 0.008 0.944 0.976
The results presented in Table 1 highlight the performance of the proposed model (including $T_{1}$ and $T_{2}$), which consistently outperforms all other approaches across the selected metrics. It achieves the lowest Mean Absolute Error, improving between 15.63% and 84.07% over other techniques; the lowest Mean Squared Error, with improvements ranging from 44.14% to 96.92%; the highest $R^{2}$ score, with gains of 4.96% to 213.54%; and the highest Pearson correlation, improving by 2.59% to 61.05%. The results including $T_{1,2}$ are a significant improvement over all models, demonstrating the importance of incorporating coherence noise into the depolarizing model. This suggests that including additional noise factors can significantly enhance the model’s accuracy in predicting the fidelity of quantum circuits executed on a real device, making it a highly promising method for future quantum error modeling and prediction.
Since the QVA model already predicts a lower fidelity than the observed success rate even without accounting for $T_{1}$ and $T_{2}$ errors, incorporating coherence errors further worsens its predictions. This underscores the need for a more robust theoretical foundation.
To ensure the robustness of our results, we repeated the experiments multiple times, conducting a total of six independent runs. In all cases, the proposed model consistently achieved the best performance across different calibration instances and quantum processors. Supplementary Figure 8 provides a detailed comparison of the prediction differences observed across all estimation methods, as well as the obtained $R^{2}$ scores for each experiment.
3 Discussion
The performance differences between the considered fidelity estimation models reveal key insights into their underlying assumptions and applicability. The QVA model, despite showing competitive performance in regression metrics, predicts lower fidelities than those observed in real hardware for some circuits. This behaviour, combined with its exclusion of significant error sources such as coherence noise, suggests incomplete theoretical foundations. While the QVA model may provide reasonable predictions in some contexts, its tendency to deviate from hardware-observed outcomes raises questions about its reliability in diverse scenarios. It is, therefore, crucial not only to evaluate models based on final metrics but also to examine their assumptions and mechanisms, ensuring they align with the complexities of real devices.
Unlike simulations requiring extensive computational resources, the scalable depolarizing model proposed in this work can efficiently and accurately predict circuit fidelities, opening new avenues for exploration. For instance, it enables comparative studies of compilation strategies seeking to maximize computation fidelity, as showcased in existing literature [30]. Additionally, the model facilitates the exploration of computation bounds for varying architectural parameters. By adjusting single- and two-qubit depolarizing factors ($p_{1}$ and $p_{2}$), the model can determine the conditions necessary to achieve high-fidelity computation. Figure 5 serves as an example. It illustrates, for various quantum circuits, the required $p_{2}$ values for achieving a circuit fidelity above 0.99 as the number of qubits increases for fixed $p_{1}$ values.
These types of analysis are only possible thanks to the proposed estimation method’s accuracy and scalability. If the model were inaccurate, it would lead to misleading predictions, potentially guiding hardware design and algorithmic decisions in the wrong direction. Additionally, without it being scalable, it would be unfeasible to thoroughly explore the full range of architectural parameters or extend the analysis to a larger number of qubits. By enabling precise and efficient fidelity estimations, the proposed approach allows for an in-depth evaluation of different quantum computing scenarios, supporting the development of more reliable and optimized quantum hardware and software.

### v2

# An Accurate and Efficient Analytic Model of Fidelity Under Depolarizing Noise Oriented to Large Scale Quantum System Design

@@ Abstract
Fidelity is one of the most valuable and commonly used metrics for assessing the performance of quantum circuits on error-prone quantum processors. Several approaches have been proposed to estimate circuit fidelity without executing it on quantum hardware, but they often face limitations in scalability or accuracy. In this work, we present a comprehensive theoretical framework to predict the fidelity of quantum circuits under depolarizing noise. Building on theoretical results, we propose an efficient fidelity estimation algorithm based on device calibration data. The method is thoroughly validated through simulation and execution on real hardware, demonstrating improved accuracy compared to state-of-the-art alternatives, with enhancements in prediction $R^{2}$ ranging from 4.96% to 213.54%.. The proposed approach provides a scalable and practical tool for benchmarking quantum hardware, comparing quantum software techniques such as compilation methods, obtaining computation bounds for quantum systems, and guiding hardware design decisions, making it a critical resource for developing and evaluating quantum computing technologies.
@@ 1 Introduction
Assessing the performance of quantum processors is a fundamental step in advancing quantum computing [2]. Reliable performance evaluation enables the design of future quantum architectures, facilitates the comparison of quantum software tools such as compilation strategies, and supports informed decision-making at all stages of quantum system development. As quantum hardware scales, understanding the impact of noise and errors on computations becomes crucial for optimizing both hardware and software components for practical applications [1].
Several methods have been proposed to estimate the reliability of quantum computations [20, 21, 22, 23, 24, 25], offering insights into the expected performance of quantum algorithms. However, existing approaches typically suffer from one of three significant limitations. Some methods lack scalability, restricting their applicability to small-scale quantum processors and making them impractical for analyzing larger, more complex systems [20]. Others lack a strong theoretical foundation, resulting in predictions that fail to reflect the behaviour of the quantum states accurately [21, 22]. Additionally, many approaches struggle to generalize across different hardware platforms, requiring adaptation or retraining when used beyond their originally targeted systems [23, 24, 25]. These limitations hinder the ability to make precise performance assessments, which are essential for benchmarking hardware, refining error mitigation techniques, and guiding quantum system design.
This work proposes a theoretical model that analytically characterizes how depolarizing noise affects quantum states when executing a circuit on a quantum processor. This model serves as the foundation for a fidelity estimation algorithm that predicts the final quantum state fidelity for a given quantum circuit and calibration data of the quantum processor.
The rest of the paper is structured as follows: Section 2 presents the necessary background for this work and reviews state-of-the-art fidelity estimation techniques. Section 3 studies the effects of depolarizing noise on quantum states and develops the theoretical framework underlying our fidelity estimation algorithm. Section 4 details the experimental methodology, and Section 5 presents and discusses the results obtained. Section 6 explores potential use cases for the proposed model and its broader implications. Finally, Section 7 summarizes the key findings and conclusions of this study.
2 Background and Related Work
Quantum fidelity [26, 27] is a measure of the similarity between two quantum states and serves as an indicator of either the reliability of a computation or the presence of excessive errors. It is calculated using the density matrix representation of the quantum states ($\rho$ and $\sigma$), as defined in Equation (1).
@@ 2 Background and Related Work
Estimating the fidelity of quantum circuits without executing them on quantum hardware is crucial to advancing the design of quantum processors. A reliable fidelity estimation method enables researchers to compare quantum systems, optimize compilation strategies, and predict performance across different hardware platforms. As quantum circuits grow in complexity with increasing qubit numbers, the estimation method must stay computationally efficient to ensure scalability.
@@
When simulating quantum circuits [28, 29, 30, 31] noise can be added to the quantum state, a process known as statistical fault injection [20]. The fidelity of the computation can be assessed by comparing a simulated noisy state with its noiseless counterpart. However, the simulation of quantum states is computationally expensive, with complexity growing exponentially with the number of qubits, restricting its application to small-scale circuits [32]. The accuracy of the estimation will depend on the noise model used in the simulation, as well as on the precision of the simulation itself, which may introduce errors arising from approximations (e.g., limiting the dimensionality of tensors in a tensor network simulation [33]).
Analytical approaches for fidelity estimation [21, 22] typically rely on gate error rates obtained from device calibration, usually through randomized benchmarking [34, 35], without assuming any error model in particular. An example of such methods is the Probability of Successful Trial (PST) [21], also known as Estimated Success Probability (ESP), a commonly used metric for estimating the fidelity of a quantum computation by multiplying the individual gate fidelities ($F_{g}$) and measurement fidelities ($F_{m}$) of the given circuit (Equation (4)). However, ESP does not account for how errors propagate through the system or specify which qubits are affected by the errors, obtaining a circuit fidelity that does not depend on the circuit’s structure but just on the number of gates and their respective fidelities.
@@
The Quantum Vulnerability Analysis (QVA) method [22], like ESP, uses individual gate fidelities as its basis. However, it introduces a hyperparameter ($w\in[0,1]$) that quantifies the degree of cross-error introduced by a two-qubit gate, requiring fine-tuning for each specific system. Additionally, unlike ESP, QVA accounts for the fidelity of each two-qubit gate twice (once for each qubit involved), resulting in consistently lower fidelity estimates compared to ESP.
Machine-learning-based techniques [23, 24, 25] have also been proposed to estimate circuit fidelity. These approaches often target specific platforms and require extensive training data and retraining for different devices. This dependence on platform-specific training limits their adaptability and scalability, making them less suitable for general-purpose fidelity estimation.
To better understand the strengths and limitations of existing fidelity estimation methods, we compare their predictions against experimental results from an actual quantum processor. Figure 1 presents the fidelity predictions of statistical fault injection (using Qiskit’s simulator [36]), ESP, and QVA for quantum circuits ranging from 2 to 8 qubits, alongside the success rate of the same circuits executed on IBM Q Kyiv (Eagle r3 processor [37]). The process from obtaining the circuits to computing the success rate after readout mitigation is explained in Section 4.2.
Figure 1 shows how both Qiskit’s simulation and ESP inaccurately predict the fidelity observed in real quantum processors, consistently overestimating it. Conversely, QVA’s performance highly depends on its hyperparameter $w$, with fidelity predictions varying significantly, up to 0.8, between the extreme values of $w=0$ and $w=1$ (QVA max and QVA min in Figure 1). This high variability introduces the additional challenge of determining the optimal $w$, which the original work [22] addressed by training a machine learning model for this purpose. In our analysis, we use $w=0.5$ (depicted as red points in Figure 1) as it provides a balanced prediction (see box plot in Figure 1) while also reporting results for $w=0$ and $w=1$. In the original work [22], the authors identify $w=0$ as a near-optimal value for large circuits. Therefore, we report results for both $w=0.5$ and $w=0$ when assessing fidelity estimation performance to ensure a fair comparison.
@@
These observations highlight the need for a more accurate fidelity estimation model that, like QVA and ESP, is scalable in terms of computational cost, allowing the number of qubits and gates in the circuit to increase. However, minimizing the dependence on hyperparameters that can excessively influence predictions is essential, as this variability can compromise the reliability and consistency of the model’s outcomes.
In this work, we develop a theoretical model for estimating the fidelity under depolarizing noise (Equation (5)), under the assumption that each quantum gate introduces depolarization to the involved qubits, a widely accepted perspective on operational errors [38, 39, 40, 41, 42, 2]. By examining how depolarizing noise impacts circuit fidelity, we lay the groundwork for a scalable fidelity estimation algorithm rooted in theoretical principles.
@@
We primarily use the last expression in Equation (5) to represent the depolarizing channel, where $d$ is the dimensionality of the quantum system being depolarized ($2^{n}$ for an n-qubit system), $I_{d}/d$ is the maximally mixed state with the same dimensions ($d$) as $\rho$ (the quantum state described by its density matrix), and $p$ is the depolarization factor (or depolarization probability). This formulation provides a compact and general framework to model depolarization effects across systems of varying dimensions.
In the following two sections, we present our proposed fidelity estimation methodology. Section 3 develops the theoretical foundation of our approach, deriving a closed-form analytical model that characterizes how depolarizing noise affects quantum states throughout a circuit. Building on these theoretical insights, Section 4 introduces a practical fidelity estimation algorithm that leverages the model to provide accurate predictions using only the quantum circuit structure and calibration data from the target device.
3 Theoretical Modeling of the Fidelity Loss
@@
To further explore the impact of depolarization, Corollary 1 analyses the fidelity loss after a single depolarization, considering the effects of repeated depolarizations over time. This finding serves as a basis for quantifying the contribution of individual gate errors when estimating the fidelity of quantum circuits.
@@ 3.1 Theoretical Model Validation
Figure 2 presents a comparison between the fidelities predicted by our theoretical model and those obtained from statistical fault injection simulations. The results span 93 quantum circuits, varying in size from 2 to 8 qubits and containing up to 5000 gates. Circuits are compiled to match the superconducting processor’s restrictions used in latter experiments (see Section 4.2).
@@
As expected, the fidelity predictions align closely with the simulated values, though slight discrepancies arise due to the theoretical upper and lower bounds employed in our model. To better illustrate these discrepancies, Figure 2 also shows the absolute difference between the predicted (with $p_{ent}=0.5$) and simulated fidelities. The observed deviations remain within a reasonable range (for all data points, the maximum absolute fidelity difference remains under 0.07), confirming that our approach provides a robust approximation of circuit fidelity while being significantly more efficient than direct simulation.
@@
Figure 2: Comparison of predicted versus simulated fidelities with single-qubit gates depolarization of $p_{1}=10^{-3}$ and two-qubit gates depolarization of $p_{2}=5\cdot 10^{-3}$. In the left plot, each point illustrates the predicted ($x$-axis) against the simulated fidelity ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. Horizontal lines indicate the range of predicted fidelities arising from the variability of the hyperparameter $p_{ent}$. The right plot shows the difference in fidelity ($fid_{pred}-fid_{simul}$).
The results in Figure 2 also reveal that the fidelity range determined by $p_{ent}=0$ and $p_{ent}=1$ is relatively narrow, in contrast to QVA, where the hyperparameter $w$ can cause fidelity estimates to vary by as much as 0.8 (see Figure 1). This demonstrates that an intermediate value, such as $p_{ent}=0.5$, can reasonably estimate the fidelity. Consequently, the inability to precisely determine the level of entanglement does not significantly impact the accuracy of the fidelity estimation.